OpenStudy (anonymous):

Original problem was (6,-3), m=- 4/5 write an equation for each line passing through the given point and having the given slope.give final answer in slope intercept.

OpenStudy (johnweldon1993):

so you have a point....and a slope...use the point slope formula y - y1 = m(x - x1) now plus in youe slope for m...your x for x1....and your y for y1

OpenStudy (anonymous):

-3=-24/5=b correct. Now what do I do next?

OpenStudy (anonymous):

-3=-24/5+b sorry, the other one was incorrect

OpenStudy (johnweldon1993):

hang on ..lets just plug in the numbers here y - (-3) = -4/5(x - 6) y + 3 = -4/5x + 24/5 y = -4/5x + 39/15 simplify it a bit to y = -4/5x + 13/5 and that just happens to be in slope intercept form

OpenStudy (anonymous):

y=m x + b Plug in the values of x, y and m, 6, -3 and -(4/5) respectively and then solve the result for b.

OpenStudy (anonymous):

\[y=-\frac{4 x}{5}+\frac{9}{5} \]

OpenStudy (johnweldon1993):

@robtobey is correct...i added somewhere where i should have subtracted